# Data-Driven Dynamic Factor Modeling via Manifold Learning

### v1

@@ Appendix A Diffusion Kalman Filter
Assuming that the gradients in (3.8) are (approximately) constant, [shnitzer2020diffusion] define a state-space model for the diffusion coordinates using a Kalman filter ($x_{t}$ is the measurement and $\psi_{t}$ diffusion coordinates):
@@
${-\varepsilon^{-1}\log\kappa_{1},\dots-\varepsilon^{-1}\log\kappa_{\ell}}$ on its diagonal, and $\mathbf{H}$ is the lifting operator. Note that ${w}$ and ${v}$ are Gaussian noises with covariance matrices $\mathbf{Q}$ and $\mathbf{V}$ respectively. Denote $\hat{\psi}_{t}=\mathbb{E}[\psi_{t}|\mathcal{F}_{t-1}]$, where $\mathcal{F}_{t-1}=\left\{x_{0},\ldots,x_{t-1}\right\}$, and $P_{t}=\operatorname{Cov}(\psi_{t}|\mathcal{F}_{t-1})$
@@
Appendix B Joint Diffusion Kalman Filter Explanatory Power
@@
Figure B.1: Plot of the $R^{2}$ from the individual stock return time series regressions directly from the JDKF state-space fitted observation equation
Figure B.2: Plot of the $R^{2}$ from the OLS Regressions of Individual Stocks vs Original Set of 132 Macro Variables
Figure B.3: Plot of the $R^{2}$ from OLS Regressions of Individual Stocks vs Reduced Set of 39 DKF Diffusion Coordinates
Figure B.4: Plot of the $R^{2}$ from OLS Regressions of Individual Stocks vs Reduced Set of 39 JDKF Diffusion Coordinates
Appendix C Second Moment Computation
We provide the full computation for the second moment of $V$ from the proof of Theorem 4.3. A Taylor approximation following [singer2008nonlinear, Equation (28)] yields
@@ Appendix C Second Moment Computation
where $E$ is a scalar function (see [singer2008nonlinear, Paragraph following Equation (28)]) which will cancel out in the following computation. In what follows, we will suppress dependence on $w$, and all expectations will be taken with respect to the invariant measure $\mu$.
@@
[…]
@@
[…]
@@
$=p^{2}(\Delta(pf^{2})-f^{2}\Delta p-2f(\Delta(pf)-f\Delta p))$
So that
[…]
@@
Appendix D Dynamic PCA Kalman Filter Equations
We provide the standard EM algorithm procedure and Kalman filter equations for the Dynamic PCA benchmark outlined in Section 5.1.3 below:
D.1 E-Step: Kalman Filter with Joint Observations
The Kalman filter is used to estimate the latent state $\mathbf{Z}_{t}$ and the corresponding state covariance $P_{t|t}$. Note that for Dynamic PCA we utilize the standard Kalman filter and EM algorithm formulations where we only observe a single set of measurements $\mathbf{X_{t}}$ The transition matrix $\mathbf{A}$ is known and estimated by fitting a VAR(1) process to the empirical PC components computed from the data, and the state covariance matrix $\mathbf{Q}$ is computed as the empirical covariance matrix of these PC components. The observation equation is:
@@ D.1 E-Step: Kalman Filter with Joint Observations
D.2 M-Step: Parameter Updates
@@
D.3 EM Algorithm
Algorithm 7 EM Algorithm Iteration
1:Initialize the parameter $\mathbf{R}$.
2:Set tolerance $\varepsilon$ and maximum iterations $N_{\text{max}}$.
4:while not converged and $k<N_{\text{max}}$ do
5:     E-step: Run Joint Diffusion Kalman filter using current estimate $\mathbf{R}$.
7:     M-step: Update $\mathbf{R}$ using the smoothed estimates of $\mathbf{Z}_{t}$.
8:     Convergence check:
9:     if  $|\mathcal{L}^{(k+1)}-\mathcal{L}^{(k)}|<\varepsilon$ then
10:         Converged: Stop the algorithm.
11:     else
12:         $k\leftarrow k+1$
13:     end if
14:end while
Appendix E Macroeconomic Factors Data
@@
We use the macroeconomic factors from FRED-MD by [mccracken2016fred] plus the eight factors found by [goyal2008comprehensive]. We use the version of this data that was curated in [chen2020internet] and the following transformations were applied:
Table E.1: Transformations (tCodes) for Macroeconomic Variables

### v2

@@ Appendix A Diffusion Kalman Filter
Assuming that the gradients in (3.8) are (approximately) constant, [36] define a state-space model for the diffusion coordinates using a Kalman filter ($x_{t}$ is the measurement and $\psi_{t}$ diffusion coordinates):
@@
$\varepsilon^{-1}\log\kappa_{1},\dots-\varepsilon^{-1}\log\kappa_{\ell}$} on its diagonal, and $\mathbf{H}$ is the lifting operator. Note that w{w} and v{v} are Gaussian noises with covariance matrices $\mathbf{Q}$ and $\mathbf{V}$ respectively. Denote $\hat{\psi}_{t}=\mathbb{E}[\psi_{t}|\mathcal{F}_{t-1}]$, where $\mathcal{F}_{t-1}=\left\{x_{0},\ldots,x_{t-1}\right\}$, and $P_{t}=\operatorname{Cov}(\psi_{t}|\mathcal{F}_{t-1})$
@@
Appendix B Kalman Filter and EM Algorithm Details
B.1 State-Space Equations for JDKF
We provide the complete Kalman filter equations for the Joint Diffusion Kalman Filter state-space model:
State equation: $\psi_{t+1}=\mathbf{A}\psi_{t}+w_{t},\quad w_{t}\sim\mathcal{N}(0,\mathbf{Q})$ (B.1)
Observation equation: $z_{t}=\mathbf{H}^{z}\psi_{t}+v_{t},\quad v_{t}\sim\mathcal{N}(0,\mathbf{R})$
where $z_{t}=[x_{t};y_{t}]$, $\mathbf{H}^{z}=[\mathbf{H}^{x};\mathbf{H}^{y}]$, and $\mathbf{R}=\begin{bmatrix}\mathbf{R}_{x}&\mathbf{R}_{xy}\\ \mathbf{R}^{\top}_{xy}&\mathbf{R}_{y}\end{bmatrix}$.
We begin with the forward pass (Kalman Filter). Initialize with $\hat{\psi}_{0|0}$ and $P_{0|0}$. For $t=1,\ldots,T$, we have the followins steps:
B.2 E-Step: Kalman Filter with Joint Observations
  1. 1.
    Prediction Step:
    $\hat{\psi}_{t|t-1}$ $=\mathbf{A}\hat{\psi}_{t-1|t-1}$
    $P_{t|t-1}$ $=\mathbf{A}P_{t-1|t-1}\mathbf{A}^{\top}+\mathbf{Q}$
  2. 2.
    Innovation:
    $e_{t}$ $=z_{t}-\mathbf{H}^{z}\hat{\psi}_{t|t-1}$
    $S_{t}$ $=\mathbf{H}^{z}P_{t|t-1}{\mathbf{H}^{z}}^{\top}+\mathbf{R}$
  3. 3.
    Kalman Gain:
    $K_{t}=P_{t|t-1}{\mathbf{H}^{z}}^{\top}S_{t}^{-1}$
  4. 4.
    Update Step:
    $\hat{\psi}_{t|t}$ $=\hat{\psi}_{t|t-1}+K_{t}e_{t}$
    $P_{t|t}$ $=(I-K_{t}\mathbf{H}^{z})P_{t|t-1}$
  5. 5.
    Log-Likelihood Contribution:
    $\mathcal{L}^{(t)}=-\frac{1}{2}\left(\log|S_{t}|+e_{t}^{\top}S_{t}^{-1}e_{t}+(m+n)\log 2\pi\right)$
Next, we proceed with the backward pass (RTS Smoother). Initialize at $T$ with $\hat{\psi}_{T|T}$ and $P_{T|T}$ from the forward pass. For $t=T-1,\ldots,1$, we have
  1. 1.
    Smoother Gain:
    $J_{t}=P_{t|t}\mathbf{A}^{\top}P_{t+1|t}^{-1}$
  2. 2.
    Smoothed State Estimate:
  3. 3.
    Smoothed Covariance:
    $P_{t|T}=P_{t|t}+J_{t}(P_{t+1|T}-P_{t+1|t})J_{t}^{\top}$
B.3 M-Step: Parameter Updates
Given smoothed state estimates $\{\hat{\psi}_{t|T}\}_{t=1}^{T}$ from the E-step, the M-step updates are as follows:
  1. 1.
    Update $\mathbf{R}_{x}$:
    $\mathbf{R}_{x}=\frac{1}{T}\sum_{t=1}^{T}\left(x_{t}-\mathbf{H}^{x}\hat{\psi}_{t|T}\right)\left(x_{t}-\mathbf{H}^{x}\hat{\psi}_{t|T}\right)^{\top}$
  2. 2.
    Update $\mathbf{R}_{y}$:
    $\mathbf{R}_{y}=\frac{1}{T}\sum_{t=1}^{T}\left(y_{t}-\mathbf{H}^{y}\hat{\psi}_{t|T}\right)\left(y_{t}-\mathbf{H}^{y}\hat{\psi}_{t|T}\right)^{\top}$
  3. 3.
    Update $\mathbf{R}_{xy}$:
    $\mathbf{R}_{xy}=\frac{1}{T}\sum_{t=1}^{T}\left(x_{t}-\mathbf{H}^{x}\hat{\psi}_{t|T}\right)\left(y_{t}-\mathbf{H}^{y}\hat{\psi}_{t|T}\right)^{\top}$
Consider the linear Multi-Factor Case. In this case, $\mathbf{H}^{y}=\mathbf{B}\mathbf{H}^{x}$, and we additionally update:
$\mathbf{B}=\left(\sum_{t=1}^{T}y_{t}(\mathbf{H}^{x}\hat{\psi}_{t|T})^{\top}\right)\left(\sum_{t=1}^{T}\mathbf{H}^{x}\hat{\psi}_{t|T}(\mathbf{H}^{x}\hat{\psi}_{t|T})^{\top}\right)^{-1}$
Appendix C Joint Diffusion Kalman Filter Explanatory Power
@@
Figure C.1: $R^{2}$ from regressing 132 individual stock returns on JDKF diffusion coordinates ($\ell=39$) via the fitted equation $y_{t}=\mathbf{B}\mathbf{H}^{x}\psi_{t}$. Median $R^{2}=0.52$, demonstrating that dimensionality reduction from 132 macroeconomic factors to 39 coordinates retains substantial predictive information. This contrasts with standard DKF (Figures B.2–B.4), which loses explanatory power by not jointly modeling covariates and responses.
Figure C.2: Plot of the $R^{2}$ from the OLS Regressions of Individual Stocks vs Original Set of 132 Macro Variables
Figure C.3: Plot of the $R^{2}$ from OLS Regressions of Individual Stocks vs Reduced Set of 39 DKF Diffusion Coordinates
Figure C.4: Results from regressing the individual stock returns against 39 JDKF diffusion coordinates: $y_{j,t}=\beta_{j}^{\top}\hat{\psi}_{t,\textrm{JDKF}}+\varepsilon_{j,t}$. Left: $R^{2}$ values range from 0.35 to 0.65, establishing the baseline explanatory power before dimensionality reduction. Center: F-statistics are uniformly high, indicating strong joint significance of the factor set. Right: All p-values are near zero, confirming statistical significance. These results provide a benchmark for evaluating how much explanatory power is retained by reduced-dimension embeddings in Figures B.2–B.4.
Appendix D Second Moment Computation
We provide the full computation for the second moment of $V$ from the proof of Theorem 4.3. A Taylor approximation following [39, Equation (28) ] yields
@@ Appendix D Second Moment Computation
where $E$ is a scalar function (see [39, Paragraph following Equation (28) ]) which will cancel out in the following computation. In what follows, we will suppress dependence on $w$, and all expectations will be taken with respect to the invariant measure $\mu$.
@@
$\left.\quad+\{pf+\frac{\varepsilon}{2}H(pf)\}^{2}\{p+\frac{\varepsilon}{4}H(p)\}\right)+\textrm{h.o.t.}$
@@
$\left.\quad+\{2p^{2}f\frac{\varepsilon}{2}H(pf)+p^{2}f^{2}\frac{\varepsilon}{4}H(p)\}\right)+\textrm{h.o.t.}$
$=2^{d}(\pi\varepsilon)^{3d/2}\left(\frac{\varepsilon}{4}p^{2}H(pf^{2})+(\varepsilon-\varepsilon+\frac{\varepsilon}{4})p^{2}f^{2}H(p)+(-\varepsilon-\frac{\varepsilon}{2}+\varepsilon)p^{2}fH(pf)\right)+\textrm{h.o.t.}$
$=2^{d}(\pi\varepsilon)^{3d/2}\frac{\varepsilon}{4}\left(p^{2}H(pf^{2})+p^{2}f^{2}H(p)-2p^{2}fH(pf)\right)+\textrm{h.o.t.}$
@@
$=p^{2}(\Delta(pf^{2})-f^{2}\Delta p-2f(\Delta(pf)-f\Delta p)).$
Hence,
$\mathbb{E}[Y_{j}^{2}]=2^{d}(\pi\varepsilon)^{3d/2}\frac{\varepsilon}{4}p^{2}(\Delta(pf^{2})-2f\Delta(pf)+f^{2}\Delta p)+\textrm{h.o.t.}$
@@
Appendix E Dynamic PCA Kalman Filter Equations
We provide the standard EM algorithm procedure and Kalman filter equations for the Dynamic PCA benchmark outlined in Section 6.1.
E.1 E-Step: Kalman Filter with Joint Observations
The Kalman filter is used to estimate the latent state $\mathbf{Z}_{t}$ and the corresponding state covariance $P_{t|t}$. Note that for Dynamic PCA we utilize the standard Kalman filter and EM algorithm formulations where we only observe a single set of measurements $\mathbf{X_{t}}$ . The transition matrix $\mathbf{A}$ is known and estimated by fitting a VAR(1) process to the empirical PC components computed from the data, and the state covariance matrix $\mathbf{Q}$ is computed as the empirical covariance matrix of these PC components. The observation equation is:
@@ E.1 E-Step: Kalman Filter with Joint Observations
E.2 M-Step: Parameter Updates
@@
Appendix F Implementation Details for Benchmark Methods
F.1 Standard Scenario Analysis (SSA)
Under the standard multi-factor model for stock returns, if $\mathbf{X_{t}}\in\mathbb{R}^{p}$ is the $p$-dimensional vector of common factors, we define $\mathcal{S}$ to be the set containing the indices of the factors in a scenario (the factors we intend to stress), and thus $\mathcal{S}^{C}$ is the index set of those factors we leave un-stressed and which are not in our scenario. SSA then stresses the components of $\mathbf{X}_{\mathcal{S},t}$ according to a given scenario (e.g. +20+20 on the S&P index, 2-2 on the CPI index, and +5+5 on the US Dollar/Euro exchange rate), and keeps the components in $\mathbf{X}_{\mathcal{S}^{C},t}$ unchanged (i.e. equal to their current value). The new portfolio P&L or overall return $V_{t}$ is then computed with $\mathbf{Y}_{t}$ determined by the scenario and the multi-factor model. By leaving $\mathbf{X}_{\mathcal{S},t}$ unchanged, SSA implicitly assumes that $\mathbb{E}[\mathbf{X}_{\mathcal{S}^{C},t+1}|\mathcal{F}_{t},\mathbf{X}_{\mathcal{S},t+1}]=0$, here $\mathcal{F}_{t}$ denotes a filtration with information up to time $t$.
  1. 1.
  2. 2.
  3. 3.
    Obtain the predicted factor vector under SSA
  4. 4.
We summarize the SSA procedure in the following algorithm, which we will then feed into a historical rolling window backtest in the next section. We let $s$ denote the size of the rolling window. Note that the fitted matrix of factor loadings, $\mathbf{\tilde{B}}$ is treated as an input in this algorithmic format. We denote with $x_{t-s:t}\in\mathbb{R}^{s\times p}$ the matrix of common factors from time tst-s up to $t$, and with $y_{t-s:t}\in\mathbb{R}^{s\times n}$ the matrix of asset returns from time tst-s up to $t$. We take $x_{\mathrm{actual},\mathcal{S},t+1}$ to be the realized historical scenario.
Algorithm 2 SSA Portfolio Simulation
[…]
4:Compute the $1/N$ (or Markowitz) weights $w$
F.2 Static PCA
Given a factor matrix $\mathbf{X}\in\mathbb{R}^{T\times p}$ and a return matrix $\mathbf{Y}\in\mathbb{R}^{T\times n}$, the static PCA procedure comprises a series of steps. First, we estimate the linear multi-factor regression model, then we project the factor changes onto principal components, apply stress scenarios, and predict the corresponding returns from the fitted factor model.
  1. 1.
    Preprocessing and Centering.
    $\Delta\mathbf{X}_{t}=\mathbf{X}_{t}-\mathbf{X}_{t-1}.$
    $\tilde{\mathbf{X}}_{t}=\Delta\mathbf{X}_{t}-\mu_{\Delta\mathbf{X}}.$
  2. 2.
    Estimation of the Multivariate Regression Model.
    We fit the multivariate regression model to obtain $\mathbf{\tilde{B}}$:
    $\mathbf{Y}_{t}=\mathbf{X}_{t}\mathbf{B}+\mathbf{\varepsilon}_{t},$
  3. 3.
    Principal Component Analysis (PCA).
    $\mathbf{X}_{\text{PCA}}=U\Sigma V^{T}.$
    and extract the principal component loading matrix:
    $W=V^{T}.$
  4. 4.
    Applying Stressed Scenarios in PCA Space.
    $\Delta\mathbf{X}=\mathbf{X}_{t+1}^{\text{scenario}}-\mathbf{X}_{t}.$
    Center the perturbation vector:
    ${\Delta\tilde{\mathbf{X}}}=\Delta\mathbf{X}-\mu_{\Delta\mathbf{X}}.$
    Project the stress perturbation onto the PCA component space:
[…]
    Obtain the perturbed scenario factor vector:
  5. 5.
    Prediction of Stressed Returns.
We summarize the static PCA procedure in the following algorithm, which we will feed into our historical rolling window backtest in the next section. We let $s$ denote the size of the rolling window. Note that the fitted matrix of factor loadings, $\mathbf{\tilde{B}}$ is treated as an input in this algorithmic format. We denote with $x_{t-s:t}\in\mathbb{R}^{s\times p}$ the matrix of common factors from time tst-s up to $t$, and with $y_{t-s:t}\in\mathbb{R}^{s\times n}$ the matrix of asset returns from time tst-s up to $t$. We take $x_{\mathrm{actual},\mathcal{S},t+1}$ to be the realized historical scenario. For notational brevity, in the algorithm below we also assume that the mean of the factor changes, $\mu_{\Delta\mathbf{X}}$, is already added back into $\text{proj}_{\text{PCA}}$
Algorithm 3 Static PCA Portfolio Simulation
[…]
2:Apply PCA on the (pre-processed) differenced data $\Delta x_{t-s:t}$
4:Obtain the change vector $\Delta x_{t}=\tilde{x}_{t+1}-x_{t}$
8:Compute the $1/N$ (or Markowitz) weights $w$
F.3 Dynamic PCA
The procedure consists of the following steps:
  1. 1.
  2. 2.
    From the state space representation we then implement the standard Kalman filter (see Appendix E) to obtain our dynamic estimates $\mathbf{Z}_{t}$ of the PC embeddings. In our empirical experiments we select the dimension of the PC embeddings as $d=5$ such that $99\%$ of the variance in our data is explained by them.
  3. 3.
[…]
    We obtain $K$ Monte Carlo conditional samples:
    $\{\mathbf{Z}^{(k)}_{t+1}\mid\mathbf{Z}_{t},\mathbf{X}_{\mathcal{S},t}\}_{k=1}^{K}$
  4. 4.
  5. 5.
We summarize the dynamic PCA procedure post-Kalman filter fitting, in the following algorithm, which we will feed into a historical rolling window backtest in the next section. We let $s$ denote the size of the rolling window. Note that the fitted matrix of factor loadings, $\mathbf{\tilde{B}}$, the matrix of eigenvectors $\mathbf{\Gamma}$, and the time $t$ Kalman filter estimate $z_{t}$ for the latent state are all treated as inputs in this algorithmic format. We denote with $x_{t-s:t}\in\mathbb{R}^{s\times p}$ the matrix of common factors from time tst-s up to $t$, and $y_{t-s:t}\in\mathbb{R}^{s\times n}$ the matrix of asset returns from time tst-s up to $t$.
Algorithm 4 Dynamic PCA Portfolio Simulation
[…]
2:for $k\leftarrow 1$ to $K$ do
3:  Generate $z_{t+1}^{(k)}\mid(z_{t},x_{\mathcal{S},t+1})$
5:  Compute the $1/N$ (or Markowitz) weights $w^{(k)}$
7:end for
F.4 Joint Diffusion Kalman Filter (JDKF)
We describe the implementation details of the JDKF algorithm:
  1. 1.
  2. 2.
    From the state space equations, we then estimate the JDKF parameters, which include $\mathbf{B}$, $\mathbf{R}_{X}$, $\mathbf{R}_{Y}$, and $\mathbf{R}_{XY}$ by fitting the joint Kalman filter as outlined in Section 5 to obtain filtered dynamic diffusion coordinate estimates $\psi_{t}^{\mathrm{JDKF}}$, from which we also obtain $\mathbf{X}_{t}^{\mathrm{JDKF}}=\mathbf{H}_{X}\psi_{t}^{\mathrm{JDKF}}$
  3. 3.
    $\mu_{\mathrm{cond,diff}}=\mathbf{H}_{\mathcal{S}^{C}}^{+}\mu_{\mathrm{cond}}$
    We obtain $K$ Monte Carlo conditional samples:
    $\{\psi^{(k)}_{t+1}\mid\psi_{t},\mathbf{X}_{\mathcal{S},t}\}_{k=1}^{K}$
  4. 4.
    $\mathbf{Y}_{t+1,\mathrm{JDKF}}^{(k)}=\mathbf{B}\mathbf{H}_{X}\psi_{t+1}^{(k)}$
  5. 5.
We summarize the JDKF procedure post-Kalman filter fitting, in the following algorithm, which we will feed into a historical rolling window backtest in the next section. We let $s$ denote the size of the rolling window. Note that the fitted matrix of factor loadings, $\mathbf{B}$, the linear lifting operator $\mathbf{H}_{X}$, and the time $t$ Kalman filter estimate $\psi_{t}$ for the latent state are all treated as inputs to the algorithm. We denote with $x_{t-s:t}\in\mathbb{R}^{s\times p}$ the matrix of common factors from time tst-s up to $t$, and $y_{t-s:t}\in\mathbb{R}^{s\times n}$ the matrix of asset returns from time tst-s up to $t$.
Algorithm 5 JDKF Portfolio Simulation
[…]
2:for $k\leftarrow 1$ to $K$ do
3:  Generate $\hat{\psi}_{t+1}^{(k)}\mid(\hat{\psi}_{t},x_{\mathcal{S},t+1})$
5:  Compute the $1/N$ (or Markowitz) weights $w^{(k)}$
7:end for
F.5 Historical Backtesting Algorithm
The complete historical rolling backtest procedure with periodic Kalman filter refitting is as follows:
1:$s,T,K,R$
$\triangleright$ s=#s=\# periods in rolling window for model training
$\triangleright$ T=#T=\# periods in the data
$\triangleright$ K=#K=\# Monte-Carlo samples used to estimate factor model-based scenario Portfolio Return
$\triangleright$ R=#R=\# months before re-fitting the Kalman filter
$\triangleright$ $r$ is the start index for refitting
$\triangleright$ $r_{\text{end}}$ is the end index for refitting
2:for $r\leftarrow 0:R:T-1$ do
3:  $r_{\text{end}}=\min(r+R+s,T)$
6:  for $t\leftarrow\max(s,r+s)$ to $r_{\text{end}}$ do
7:   Retrieve Precomputed Kalman Filter Estimates
8:   Retrieve $\hat{\psi}_{t}$, $\hat{z}_{t}$ from Kalman filter estimates
9:   Estimate Multivariate Regression Model
10:   Fit $\mathbf{\tilde{B}}$ using $(y_{t-s:t},x_{t-s:t})$
11:   Compute Portfolio Simulations
[…]
16:   Compute true portfolio return $V_{t+1,\mathrm{TRUE}}$
17:   Compute Errors
22:  end for
23:end for
In all of our experiments in Section 7.2, we use $K=10,000$ Monte Carlo simulations of conditional portfolio returns for the probabilistic models JDKF and Dynamic PCA. Since we are working under the special linear multi-factor model case of our JDKF framework, for each refitting period we estimate $\mathbf{B}$ via ordinary least squares (OLS) and $\mathbf{R}$ as the empirical covariance of the OLS residuals; furthermore, we estimate $\mathbf{V}$ as the empirical covariance matrix of the common factors.
Appendix G Macroeconomic Factors Data
@@
We use the macroeconomic factors from FRED-MD by [30] plus the eight factors found by [21]. We use the version of this data that was curated in [10] and the following transformations were applied:
Table G.1: Transformations (tCodes) for Macroeconomic Variables